OpenStudy (anonymous):

MEDAL IF HELPED MEDAL IF HELPED MEDAL IF HELPED Solve -2x2 - 16x - 44 = 0. A. x equals negative 8 plus or minus I square root of 6 B. x equals negative 8 plus or minus 2i square root of 6 C. x equals negative 4 plus or minus i square root of 6 D. x equals negative 4 plus or minus 2i square root of 6

OpenStudy (mathstudent55):

First, divide both sides by -2. What do you get?

OpenStudy (anonymous):

x^2+8x+22=0

OpenStudy (mathstudent55):

Great. This is not factorable because there are no two numbers that multiply to 22 and add to 8.

OpenStudy (mathstudent55):

That means we use the quadratic formula.

OpenStudy (anonymous):

uhhggg i hate that word

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (mathstudent55):

It's not that bad. In your case, a = 1, b = 8, and c = 22.

OpenStudy (mathstudent55):

Now we plug in.

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) \(x = \dfrac{-8 \pm \sqrt{8^2 - 4(1)(22)}}{2(1)}\)

OpenStudy (anonymous):